OpenStudy (anonymous):

Help? Sorry for all these questions lately... Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. Part B: Write the equation obtained in Part A using function notation. Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (therealmeeeee):

y=225-2x y=480-5x and solved for y got y=55

OpenStudy (therealmeeeee):

do u got it???????????????????

OpenStudy (anonymous):

Yes, but what part is this for? sorry...

OpenStudy (therealmeeeee):

the first one

OpenStudy (therealmeeeee):

medal now

OpenStudy (anonymous):

Dang ok.